how many litres of a 40% concentrated solution and a 25% concentrated solution must be mixed to obtain 79 litres of a 32% solution?
Answer by mananth(16946)

percent ----------------quantity
40.00% ----------------x
25.00% ----------------79-x
32.00% ----------------79
....
40x+25(79-x)=32 *79

40x+1975-25x=2528

15x=553
/15
x=36.87 liters of 40%
42.13 liters of 25%